U(x, y) = x1ax2(1-a)

a. Solve for the marshallian demands for x1 and x2, as functions of p1, p2, and m. (Hint: your solutions will be equations, not numbers).

b. For x1 find the own-price elasticity and income elasticity.

c. Suppose a = 0.2, m = 100, p1 = 2, and p2=8, find the quantities of x1 and x2.

d. happens to these quantities when p1 doubles to $4?

e. What does this say about the price consumption curve (PCC)?
Answer #1

a) v(x,y) = x, ter At optimality MRS=/ Mo, P Mu-R 5) xa x +2 -1 = (2x, = h Budgetlie Bx, tht = m = PX (2) thetem &.= m(ind)